A small town produces good quality maize. Let us consider a market for the maize produced here. The farmers in this town distribute and sell their produce to three retail stores in the nearby city: Company A, Company B and Company C.

 The market supply for maize is given as Qs = 14 000 + 17.5p.  The demand curves for maize by Companies A, B and C are respectively: Qd = 40 000 – 12.5p; Qd = 30 000 – 5p; and Qd = 20 000 – 3p.

Note: P represents the different price levels for maize. Qs is the market supply of maize by the farmers from the small town. Qd is the quantity demanded by each retail store

1.calculate the equilibrium market price levels of maize
2.Make sure the graph to illustrate the market demand and market supply of maize
3.make use of a graph to show impact of minimum price fixing above the market price for maize
1
Expert's answer
2019-08-26T09:57:25-0400

1. In equilibrium Qd = Qs.

For company A:

14 000 + 17.5p = 40 000 – 12.5p;

30p = 26000,

p = 866.67.

For company B:

14 000 + 17.5p = 30 000 – 5p.

22.5p = 16000,

p = 711.11.

For company C:

14 000 + 17.5p = 20 000 – 3p,

20.5p = 6000,

p = 292.68.

2. Market demand is illustrated as a downward-sloping curve, and market supply of maize is illustrated as upward-sloping curve. They intersect in equilibrium point.

3. The minimum price fixing above the market price for maize will create a surplus of maize, because Qd < Qs.